A 75.0 kg ice hockey goalie, originally at rest, catches a 0.150 kg hockey puck slapped at him at a velocity of 18.0 m/s. Suppose the goalie and the ice puck have an elastic collision and the puck is reflected back in the direction from which it came. What would their final velocities (in m/s) be in this case? (Assume the original direction of the ice puck toward the goalie is in the positive direction. Indicate the direction with the sign of your answer.)
Answer #1

Mass of the hockey puck = m1 = 0.15 kg

Mass of the ice hockey goalie = m2 = 75 kg

Velocity of the puck before the collision = V1 = 18 m/s

Velocity of the goalie before the collision = V2 = 0 m/s

Velocity of the puck after the collision = V3

Velocity of the goalie after the collision = V4

The collision is elastic.

Coefficient of restitution = e = 1

V4 е — и — V2

V4V 1 18 0

V4 = V3 + 18

By conservation of linear momentum,

m1V1 + m2V2 = m3V3 + m4V4

(0.15)(18) + (75)(0) = (0.15)V3 + (75)(V3 + 18)

2.7 = 0.15V3 + 75V3 + 1350

75.15V3 = -1347.3

V3 = -17.928 m/s

Negative sign indicates it's direction is opposite to that of the initial velocity of the hockey puck.

V4 = V3 = 18

V4 = -17.928 + 18

V4 = 0.072 m/s

Final velocity of the puck = -17.928 m/s

Final velocity of the goalie = 0.072 m/s